# Orthogonal projections of hypercubes

### v3

@@ B.4 Minimum inner-product error formulation
$\frac{\partial}{\partial\bm{u}}\varepsilon^{2}\left(\bm{x},\bm{x}^{\prime}\right)$
$=$ $\left(\bm{x}^{\top}\bm{x}^{\prime}-\bm{u}^{\top}\bm{x}\bm{x}^{\prime\top}\bm{u}\right)\left(-2\bm{x}\bm{x}^{\prime\top}\bm{u}\right)$
@@ C.2 Quadripolar distribution
$\bm{\varPi}\bm{B}\bm{\varPi}^{\top}=$ $\left(\bm{I}-\bm{u}_{1}\bm{u}_{1}^{\top}\right)\left(-\bm{\xi}_{1}\bm{\xi}_{1}^{\top}+\bm{\xi}_{2}\bm{\xi}_{2}^{\top}\right)\left(\bm{I}-\bm{u}_{1}\bm{u}_{1}^{\top}\right)^{\top}$
@@ C.3 Relation with Hamming distance
We define the angle $\vartheta$ between the two weighted vertices as $\vartheta\coloneqq\arccos\left(\frac{\bm{\xi}_{1}^{\top}\bm{\xi}_{2}}{N}\right)$ and consider the projected coordinates of vertex $\bm{s}$ on PC1 and PC2 with zeroth-order approximation, obtaining
@@ D.1 General bound of projected coordinates
In general, the orthogonally projected coordinates of hypercubic vertices are bounded by a circle with radius $\sqrt{N}$. This is because we have $\left|\bm{r}\left(\bm{s}\right)\right|\leq\sqrt{N}$ [Eq. (44)] for any vertex. To show this, we consider the upper bound of $\left|\bm{r}\right|^{2}$,
$\left|\bm{r}\left(\bm{s}\right)\right|^{2}=$ $\bm{r}^{\top}\left(\bm{s}\right)\bm{r}\left(\bm{s}\right)$
$=$ $\sum_{i=1}^{2}\left(\bm{u}_{i}^{\top}\bm{s}\right)^{2}$
$\leq$ $\sum_{i=1}^{N}\left(\bm{u}_{i}^{\top}\bm{s}\right)^{2}$
$=$ $\sum_{i=1}^{N}\bm{s}^{\top}\bm{u}_{i}\bm{u}_{i}^{\top}\bm{s}$
$=$ $\bm{s}^{\top}\left(\sum_{i=1}^{N}\bm{u}_{i}\bm{u}_{i}^{\top}\right)\bm{s}$
$=$ $\bm{s}^{\top}\bm{s}$
$=$ $N$ (116)

### v4

@@ B.4 Minimum inner-product error formulation
$\frac{\partial}{\partial\bm{u}}\varepsilon^{2}\left(\bm{x},\bm{x}^{\prime}\right)$ $=\left(\bm{x}^{\top}\bm{x}^{\prime}-\bm{u}^{\top}\bm{x}\bm{x}^{\prime\top}\bm{u}\right)\left(-2\bm{x}\bm{x}^{\prime\top}\bm{u}\right)$
@@ C.2 Quadripolar distribution
$\bm{\varPi}\bm{B}\bm{\varPi}^{\top}$ $=\left(\bm{I}-\grave{\bm{u}}_{1}\grave{\bm{u}}_{1}^{\top}\right)\left(-\bm{\xi}_{1}\bm{\xi}_{1}^{\top}+\bm{\xi}_{2}\bm{\xi}_{2}^{\top}\right)\left(\bm{I}-\grave{\bm{u}}_{1}\grave{\bm{u}}_{1}^{\top}\right)^{\top}$
$=\left(\bm{I}-\frac{1}{N}\bm{\xi}_{1}\bm{\xi}_{1}^{\top}\right)\left(-\bm{\xi}_{1}\bm{\xi}_{1}^{\top}+\bm{\xi}_{2}\bm{\xi}_{2}^{\top}\right)\left(\bm{I}-\frac{1}{N}\bm{\xi}_{1}\bm{\xi}_{1}^{\top}\right)^{\top}$
@@ C.3 Relation with Hamming distance
We define the angle $\vartheta$ between the two weighted vertices as $\vartheta\coloneqq\arccos\left(\frac{\bm{\xi}_{1}^{\top}\bm{\xi}_{2}}{N}\right)$, i.e., $\frac{\bm{\xi}_{1}^{\top}\bm{\xi}_{2}}{N}=\cos\left(\vartheta\right)$, and consider the projected coordinates of vertex $\bm{s}$ on PC1 and PC2 with zeroth-order approximation, obtaining
@@ D.1 General bound of projected coordinates
In general, the orthogonally projected coordinates of hypercubic vertices are bounded by a circle with radius $\sqrt{N}$. This is because we have $\left|\bm{r}\left(\bm{s}\right)\right|\leq\sqrt{N}$ [Eq. (65)] for any vertex. To show this, we consider the upper bound of $\left|\bm{r}\left(\bm{s}\right)\right|^{2}$,
$\left|\bm{r}\left(\bm{s}\right)\right|^{2}$ $=\bm{r}^{\top}\left(\bm{s}\right)\bm{r}\left(\bm{s}\right)=\sum_{i=1}^{2}\left(\bm{u}_{i}^{\top}\bm{s}\right)^{2}$
$\leq\sum_{i=1}^{N}\left(\bm{u}_{i}^{\top}\bm{s}\right)^{2}=\sum_{i=1}^{N}\bm{s}^{\top}\bm{u}_{i}\bm{u}_{i}^{\top}\bm{s}=\bm{s}^{\top}\left(\sum_{i=1}^{N}\bm{u}_{i}\bm{u}_{i}^{\top}\right)\bm{s}$
$=\bm{s}^{\top}\bm{s}=N,$ (160)
which is
$\left|\bm{r}\left(\bm{s}\right)\right|\leq\sqrt{N}.$ (161)